# MIMIR: Masked Image Modeling for Mutual Information-based Adversarial Robustness

### v2

@@ 4.4 MI Measures
We also show that MIMIR outperforms original MAE [36] and Adv-MAE [91]. To make it clearer, the MAE baseline in Table VII refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD. Adv-MAE refers to using adversarial examples but not using $I(x+\delta,z)$ in loss. Adv-MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training. Adv-MAE provides better accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task. This more difficult task further improves the performance of downstream models. We use the default learning rate (i.e., $45.0e-$ 4) of MAE [36] in experiments in Table VII. This is aligned with Adv-MAE [91].
Figure 4: The fine-tuning results with different numbers of pre-train epochs. The left column is trained on CIFAR-10, and the right column is on Tiny-ImageNet. The above is natural accuracy, and the below is adversarial accuracy under a 20-step PGD attack.
4.5 Training Epochs Ablation Study
Our experiments so far are based on 800-epoch pre-training. In Figure 4, we show the influence of different numbers of epochs. We use ViT-T and ViT-S as our baseline models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet. We fine-tune each model 50 epochs after MIMIR pre-training. Both adversarial and natural accuracy are improved with a longer training schedule. On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S. On Tiny-ImageNet, the increase is more obvious.
4.6 Fine-tuning Using Advanced Methods
Our fine-tuning thus far is based on traditional 10-step PGD adversarial training for CIFAR-10. The reason is to compare with previous works that use a 10-step PGD algorithm. Next, we consider using TRADES [96] and MART [82] to study whether MIMIR can achieve better performance with better fine-tuning methods. TRADES [96] identify a trade-off between robustness and accuracy on clean data. The authors decompose the prediction error of adversarial examples as natural and boundary errors. According to these two kinds of errors, the authors find a differentiable upper bound based on the theory of classification-calibrated loss. MART [82] finds that misclassified and correctly classified natural examples have different influences on the final robustness after adversarial training. Therefore, the authors designed MART to explicitly differentiate the misclassified and correctly classified natural examples during adversarial training.
In Table VIII, the MIMIR performance could be further improved with the two advanced adversarial training methods. Specifically, TRADES slightly improves the accuracy on clean examples but shows worse robustness compared to PGD. MART improves the adversarial accuracy when using ViT-B, but we notice a significant drop in natural accuracy.
TABLE VIII: The experimental results with different fine-tuning methods.
[…]
ViT-S PGD 88.11 56.63 56.34 53.18
TRADES 88.19 56.42 56.36 51.70
MART 80.55 50.81 50.75 39.92
ViT-B PGD 89.30 58.07 57.87 54.55
TRADES 90.91 57.73 57.58 52.79
MART 85.78 60.35 60.16 54.27
4.7 Adaptive Attack Evaluation
As the adversary is aware of the defense MIMIR, they can design adaptive attacks [76]. For example, the adversary may attack feature space [52, 65] since MIMIR trains backbone to extract features. Here, the backbone refers to the ViT model without the final classification layer, i.e., the encoder of MIMIR. In this section, we provide two adaptive attacks specifically designed against MIMIR. First, we introduce PGD-MI, which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in loss. Therefore, it is reasonable to attack the model by directly increasing the MI $I(x+\delta,z)$. Formally, we embed the following loss into the PGD algorithm:
We use the same value of $\lambda$ as in MIMIR pre-training.
Second, we introduce PGD-feature that directly attacks the feature extracted by ViT backbones. As the backbones are trained by self-supervised MIMIR pre-training (the whole model is trained when fine-tuning), we hope to evaluate the feature extracted by the backbones. The PGD-feature increases the distance between features extracted from clean and adversarial examples. We implement it using the $L_{2}$ norm and the PGD algorithm:
$\mathop{\max}\limits_{\delta\in S}l_{2}$ $(f_{e}(x),f_{e}(x+\delta)).$ (22)
Table IX shows the adaptive evaluation results for PGD-MI and PGD-feature. We use 100 steps for both to make sure the attacking algorithm converges. The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$. For ImageNet-1K, we use $\epsilon=4/255$. According to the results in Table IX, MIMIR is still robust against adaptive attacks.
4.8 Evaluation of the Loss Surface
To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot loss surfaces [48] in Figure 5. The loss surface is the visualization of the loss function as parameters change. The basic idea is to plot the loss around the optimal parameters. Formally, we consider in the 2D case:
$f_{l}(\alpha,\beta)=L(\theta^{*}+\alpha\theta_{1}+\beta\theta_{2}),$ (23)
where $\theta_{1}$ and $\theta_{2}$ are two direction vectors, $\alpha$ and $\beta$ are two arguments of $f_{l}$. In practice, we use the parameters of trained models, i.e., $\theta^{*}$. In Figure 5, we provide the loss surfaces for models. The surfaces of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
Figure 5: The loss surfaces of the models. The four figures of each column are the loss information of one model. The first column is ConViT-S pre-trained with MIMIR and fine-tuned with adversarial examples. The second column is ConViT-S, pre-trained with clean data and fine-tuned with adversarial examples. The third is ViT-S, pre-trained with clean data and then fine-tuned with adversarial examples. The fourth one is ViT-S, pre-trained with MIMIR and fine-tuned with adversarial examples. The training dataset is CIFAR-10.
TABLE IX: The experimental results for adaptive attacks.
[…]
@@ 4.8 Evaluation of the Loss Surface
4.9 Efficiency
Table X provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning. The time consumption and memory usage are evaluated on 4 A6000 GPUs. Compared to 10-step PGD and FAST AT, MIMIR shows much less time consumption. At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$. Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time. We also consider using 10-step PGD at pre-training, Table VII shows that using a one-step algorithm with larger $\alpha=10$ provides better performance. In addition, pre-training with MIMIR also costs less memory. Note in Table X that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning. In Table X, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training [86]. Therefore, MIMIR is more promising concerning training costs when conducting adversarial training.
4.10 Discussion and Limitations
Across our experiments, we have observed the promising performance of MIMIR. This is because we build a different adversarial training paradigm. With MIMIR, the adversarial pre-training can be more efficient and effective. Following the idea of IB, we can intuitively consider a bottleneck between the encoder and decoder. As the reconstruction output is constrained by clean data $x$, the bottleneck will filter out information from adversarial perturbation $\delta$. We provide a theoretical guarantee of this bottleneck in Section 3.4. In Eq. (8), we embed this bottleneck as a learning object to further improve the performance. This improvement also confirms the correctness of our theoretical guarantee in Section 3.4. Table VII shows that our method works better than baseline even without embedding the bottleneck in Eq. (8). With the two information sources of $x$ and $\delta$, the model is trained to learn the robust features from $x$ and forget the information of $\delta$ under the constraint of the reconstruction target.
While MIMIR shows better performance, there are still certain limitations. MIMIR is a pre-training method. An adversarial fine-tuning is necessary to build the final model. Thus, the shortcomings of traditional adversarial training cannot be completely avoided. In our experiments, we utilize the simple PGD algorithm for fine-tuning, but one can further improve MIMIR pre-trained models with more advanced approaches. In addition, MIMIR follows the design of MAE, and we also utilize the characteristic that ViTs are capable of processing variable-length inputs. Therefore, currently, MIMIR cannot handle pyramid-based ViTs and CNNs. While it is not trivial to apply MIMIR to other pyramid-based ViTs and CNNs, this is feasible [89, 49, 81].
TABLE X: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
CIFAR-10 [46] Tiny-ImageNet [47] ImageNet-1K [26]
Models #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
[…]
FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
[…]
FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
[…]
FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
5 Related Work
5.1 Vision Transformer
The transformers [77] were first proposed in natural language processing (NLP). With the mechanism of global self-attention, transformers can effectively capture the non-local relationships among all text tokens [27, 14, 22]. The pioneering work, ViT [28], demonstrated that the pure transformer architecture can achieve competitive performance on various tasks. ViT also reveals that transformers lack inductive biases [28]. For example, locality, two-dimensional neighborhood structure, and translation equivariance are inherent to CNNs but not applicable to ViTs [28]. Due to this shortcoming, ViTs usually require large-scale training to get competitive performance, such as pre-training on ImageNet-21K [26] and JFT-300M [70]. To alleviate the ViT need for large datasets, DeiT [74] introduced a teacher-student strategy to distill knowledge from a teacher CNN for a student ViT. In the MIM field, MAE [36] uses a masked autoencoder with a lightweight decoder as a visual representation learner. Its learning objective is to reconstruct the original image by the decoder while using masked images as input to the autoencoder. The advantage is that MAE can randomly discard 75% image patches when pre-training under ImageNet-1K [26], which means more efficient training.
5.2 Adversarial Attacks on ViTs
The concept of adversarial attacks first appeared in [23], which proposed a formal framework and algorithms against the adversarial spam detection domain. Then, the adversarial attacks were popularized by Biggio et al. [11] and Szegedy et al. [71] in image classification. The generation of adversarial examples depends on the model’s gradient or estimated gradient in black box situation [43]. Therefore, adversarial attacks can be easily applied to transformers by using the gradient of attention blocks concerning inputs. This also raises a question: are transformers more robust than CNNs? Benz et al. [9] found that CNNs are less robust than ViTs due to their shift-invariant property. Bhojanapalli et al. [10] found that ResNet models are more robust than transformers at the same model size under FGSM attack, but under PGD [55] attack, transformer models show better robustness. As ViTs process the input image as a sequence of patches, Gu et al. [35] found that ViTs are more robust than CNNs to naturally corrupted patches because the attention mechanism is helpful in ignoring naturally corrupted image patches. The later work [6] revealed that CNNs could be as robust as ViTs against adversarial attacks if CNNs are trained with proper hyperparameters.
5.3 Adversarial Defense
Due to the difference between CNNs and ViTs, there have been some recent efforts to explore new adversarial training approaches for ViTs [57, 25, 87]. Mo et al. [57] presented a new adversarial training strategy based on the following observations: 1) pre-training with natural data can provide better robustness after adversarial fine-tuning, 2) gradient clipping is necessary for adversarial training, and 3) using SGD as the optimizer is better than Adam. Debenedetti et al. [25] also presented an improved training strategy for ViTs by evaluating different combinations of data augmentation policies. They found that weak data augmentation and large weight decay are better than previous canonical training approaches. As adversarial training is time-consuming, AGAT [87] leverages the attention score while training to discard non-critical image patches after every layer. Unlike previous works, we provide a different training paradigm by using MIM for adversarial pre-training. Our method is efficient as we discard 75% image patches while pre-training. Our method is effective as we eliminate the information of adversarial perturbation from two information sources of clean and adversarial inputs. We also provide theoretical proof that the information of adversarial perturbation is eliminated.
5.4 Self-Supervised Adversarial Pre-Training
[…]
@@ 6 Conclusions and Future Work
This paper investigates MIMIR as a pre-training method to improve adversarial robustness for ViTs under various benchmark datasets. MIMIR uses adversarial examples as inputs and clean data as the reconstruction target. In this way, the information from the adversarial perturbation is removed by the bottleneck, but the information from clean data is maintained due to the constraint while reconstructing the target. We also theoretically prove that MI between adversarial inputs and their latent representations, i.e., $I(x+\delta,z)$, should be decreased while training. Our experimental results show that MIMIR significantly improves adversarial robustness compared to recent state-of-the-art methods. In addition, a single epoch of MIMIR pre-training costs only about $\frac{1}{8}$ of training time compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training [86] (see Table X). The aforesaid limitations of MIMIR represent a good starting point for future work. For example, we will investigate how to apply MIMIR to CNNs and pyramid-based ViTs and build new self-supervised tasks for adversarial training.
@@ Appendix A Datasets
We use three commonly used datasets to evaluate MIMIR: CIFAR-10 [46], Tiny-ImageNet [47], and ImageNet-1K [26]. CIFAR-10 [46] comprises 50,000 images with size $3\times 32\times 32$ in 10 classes. ImageNet-1K [26] is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes. In our experiments, images from ImageNet-1K are resized to $3\times 224\times 224$. For completeness, we also include Tiny-ImageNet [47] as a medium size dataset between CIFAR-10 [46] and ImageNet-1K [26]. Tiny-ImageNet [47] contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
@@ Appendix B Decoder Hyperparameters
We use transformer blocks but fewer layers as the backbone of the decoder. For CIFAR-10, we use the patch size of 2, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Table XI shows the hyperparameters of decoder architectures. For different ViT architectures, we use the transformer blocks of the respective architectures to build the decoder.
TABLE XI: Model architectures of the encoder and decoder.
@@
Appendix C Data Augmentation Ablation Study
Figure 6 demonstrates the loss and accuracy while training with different augmentations. “no mix” refers to using only weak augmentation, including RandomResizedCrop and RandomHorizontalFlip. “+mix” refers to using MixUp (0.8) and CutMix (1.0). “+aug” refers to using MixUp (0.8), CutMix (1.0), and Randaugment (rand-m9-mstd0.5-inc1).
Figure 6: The training results of using different data augmentations with 800 epochs.
Appendix D Details of Training Hyperparameters
In Tables XII and XIII, we provide the default hyperparameters used in our experiments. In addition, we use different patch sizes for different datasets: patch size 2 for CIFAR-10, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Using smaller patch sizes increases the time consumption when calculating self-attention, but MIMIR pre-training discards 75% patches, making it still efficient. In addition, due to the depth and comparatively small embedding size of CaiT, we use different drop path and layer-wise decay at fine-tuning (for ImageNet-1K). For CaiT-XXS24, we use 0.95 and 0.15 as layer-wise decay and dropout, and 0.85 and 0.35 for CaiT-S36. We also apply stochastic depth decay rule [42] to CaiT. CaiT-S36 models are only fine-tuned 50 epochs due to time-consumption and it is sufficient to get superior results. The batch size to fine-tune CaiT is 512 due to limitation of GPU memory. Other hyperparameters are consistent with Tables XII and XIII.
TABLE XII: Pre-training hyperparameters.
@@ Appendix D Details of Training Hyperparameters
TABLE XIII: Fine-tuning hyperparameters.
@@
layer-wise lr decay 0.65
@@
warmup epochs 10
training epochs 100
@@
Appendix E Fine-tuning with Clean Images
In Table XIV, we show the results of MIMIR pre-training with adversarial data, and MAE [36] pre-training with clean data. MIMIR shows comparable and even better performance than MAE when fine-tuning with clean data. The results in Table XIV is reported with 800 pre-training epochs. The blr used in Table XIV is 0.001, and we fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K.
TABLE XIV: The experimental results for fine-tuning with clean images.
Model Dataset Pre-train Natural Acc.
ViT-B CIFAR-10 MIMIR 96.91
MAE 96.79
Tiny-ImageNet MIMIR 75.43
MAE 73.38
ImageNet-1K MIMIR 83.20
MAE 82.92
ConViT-S CIFAR-10 MIMIR 95.38
MAE 94.95
Tiny-ImageNet MIMIR 70.40
MAE 69.03
ImageNet-1K MIMIR 79.21
MAE 78.37
Appendix F Dropout is Important for CaiT
We also consider applying our method to deeper ViTs (i.e., CaiT [75]). CaiT is designed for deeper high-capacity transformers that benefit from depth. Simply increasing the depth may cause training failure of ViTs [75]. The CaiT solves this problem by LayerScale and class-attention. In adversarial training, training deeper networks is even more difficult. The dropout has a great impact on robustness. In Table XV, we show the results using MIMIR with different dropout values.
TABLE XV: The experimental results on CaiT-S36. The “sdd” refers to stochastic depth decay rule [42].
Dropout Layer Decay LR $\epsilon=2/255$
Natural PGD
0.1 0.65 72.68 35.60
0.1+SDD 0.65 73.06 38.88
0.2 0.75 75.41 43.99
0.25 0.75 71.46 48.86
0.25 0.85 76.43 53.49
0.25+SDD 0.85 75.92 56.53
0.35+SDD 0.85 76.05 56.78
Appendix G Mutual Information and HSIC

### v3

@@ 4.2 Experimental Results
1-step is better than 10-step of AT in pre-training. We also show that MIMIR outperforms original MAE (He et al. 2022) and adv MAE with different PGD steps (to generate adversarial examples for training). MAE in Table 5 refers to using the original MAE for pre-training and then fine-tuning with 10-step PGD. The adv MAE refers to using adversarial examples without the MI $I(x+\delta,z)$ in loss. The adv MAE (10-steps) refers to using the 10-step PGD algorithm ($\epsilon=8,\alpha=2$) to generate adversarial examples at pre-training. The adv MAE provides higher accuracy than MAE, which supports our statement that using adversarial examples in Masked Image Modeling creates a more difficult reconstruction task. This more difficult task further improves the performance of downstream models (See also Table 7). We use the default learning rate (i.e., $45.0e-$ 4) of MAE, so there is a performance drop in experiments in Table 5 and 4 since AT prefers larger learning rates on CIFAR-10 as shown in Table 11 in the Appendix.
Stability. Due to the high computation cost, we cannot report standard deviation for all experiments. To show that our method MIMIR has low variances, we train ViT-S on CIFAR-10 three times with different random seeds, running pre-training 400 epochs and fine-tuning 50 epochs. The natural accuracy is 86.07 $\pm$ 0.16 %. The adversarial accuracy under PGD20 ($l_{\infty},\epsilon=8/255$) is 47.24 $\pm$ 0.12 %.
Table 6: Adversarial accuracy by adaptive attacks. The models are pre-trained 800 epochs by MIMIR and fine-tuned 100 epochs by PGD AT.
Dataset Model PGD20 PGD-MI100 PGD-fea100
@@
Adaptive Attacks. We evaluate MIMIR against adaptive adversaries following common practices (Tramer et al. 2020). We provide two adaptive attacks specifically designed against MIMIR. First, we introduce the PGD Mutual Information attack (PGD-MI), which utilizes the MI $I(x+\delta,z)$ to generate adversarial examples, as $I(x+\delta,z)$ is used in MIMIR pre-training as a penalty in the loss. PGD-MI attacks the model by directly increasing the MI $I(x+\delta,z)$. Specifically, we add the MI loss into the PGD algorithm:
where the value of $\lambda$ in MIMIR pre-training is available to adversaries.
Second, we introduce a PGD feature attack (PGD-fea) that directly attacks the feature extracted by ViT backbones following (Sabour et al. 2016). In particular, we attack the feature extractor from the backbones after the adversarial fine-tuning. The PGD-fea attack increases the Euclidean distance between features extracted from natural and adversarial examples. We implement it using the PGD algorithm:
$\mathop{\max}\limits_{\delta\in S}\mathcal{L}_{\text{mse}}$ $(f_{e}(x),f_{e}(x+\delta)).$ (19)
Both PGD-MI and PGD-fea are optimized for 100 steps to make sure the attacking algorithm converges. The perturbation budget is the same as the previous evaluation, i.e., $\epsilon=8/255$ for CIFAR-10 and Tiny-ImageNet, and $\epsilon=4/255$ for ImageNet-1K. Table 6 demonstrates the adaptive evaluation results for PGD-MI and PGD-fea attacks, showing that MIMIR is robust against adaptive attacks on benchmark datasets.
Table 7: Natural accuracy of MAE and MIMIR (800 epochs pre-training for both) that are fine-tuned on natural images.
Architecture Pre-train CIFAR-10 Tiny ImageNet
ViT-B MAE 96.79 73.38 82.92
MIMIR 96.91 75.43 83.20
ConViT-S MAE 94.95 69.03 78.37
MIMIR 95.38 70.40 79.21
ViT-S MAE 95.95 70.00 77.45
MIMIR 95.95 71.14 78.69
Figure 3: The loss landscapes of MIMIR pre-trained models.
Visualization of the Loss Landscape. To show that the robustness of MIMIR-trained models does not stem from gradient masking, we plot the loss landscape (Li et al. 2018) in Figure 3. The loss landscape is the visualization of the loss function as parameters change. The basic idea is to plot the loss around the optimal parameters. The landscapes of all models are smooth, i.e., the gradient at a certain point is clear and can also be easily estimated by local average gradients, which means the gradient is masked.
Fine-tuning with Natural Images. In Table 7, we show the results of MIMIR and the original MAE (He et al. 2022) with the same hyperparameter. We fine-tune 50 epochs for CIFAR-10 and Tiny-ImageNet, 100 epochs for ImageNet-1K. Surprisingly, MIMIR outperforms MAE when fine-tuning with natural data. We leave further exploration of this finding to future work.
@@ 5 Conclusions and Future Work
This paper provides a novel theoretical MI analysis on ViT adversarial training, showing that MI between the adversarial example and its latent representation in ViT-based autoencoders, i.e., $I(x+\delta,z)$, should be constrained by utilizing the MI bounds. Based on this finding, we propose MIMIR as a theoretically grounded pre-training method to improve adversarial robustness for ViTs. MIMIR uses adversarial examples as inputs and natural data as the reconstruction target. In this way, the information from the adversarial perturbations is decreased by the bottleneck, while the information from natural data is preserved while reconstructing the target. Our experimental results show that MIMIR substantially improves adversarial robustness compared to recent related works on various benchmark datasets.
@@ Appendix A Datasets
We use three commonly used datasets to evaluate MIMIR: CIFAR-10 (Krizhevsky, Hinton et al. 2009), Tiny-ImageNet (Le and Yang 2015), and ImageNet-1K (Deng et al. 2009). CIFAR-10 (Krizhevsky, Hinton et al. 2009) comprises 50,000 images with size $3\times 32\times 32$ in 10 classes. ImageNet-1K (Deng et al. 2009) is the most commonly used dataset for the evaluation of ViTs and its variants, which is composed of more than 1.2 million high-resolution images in 1,000 classes. In our experiments, images from ImageNet-1K are resized to $3\times 224\times 224$. For completeness, we also include Tiny-ImageNet (Le and Yang 2015) as a medium size dataset between CIFAR-10 (Krizhevsky, Hinton et al. 2009) and ImageNet-1K (Deng et al. 2009). Tiny-ImageNet (Le and Yang 2015) contains 100,000 images with size $3\times 64\times 64$ in 200 classes.
@@ Appendix B Decoder Hyperparameters
We use transformer blocks but fewer layers as the backbone of the decoder. For CIFAR-10, we use the patch size of 2, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Table 8 shows the hyperparameters of decoder architectures. For different ViT architectures, we use the transformer blocks of the respective architectures to build the encoder.
Table 8: Model architectures of the encoder and decoder.
@@
Appendix C Details of Training Hyperparameters
In Tables 9 and 10, we provide the default hyperparameters used in our experiments. In addition, we use different patch sizes for different datasets: patch size 2 for CIFAR-10, 4 for Tiny-ImageNet, and 16 for ImageNet-1K. Using smaller patch sizes increases the time consumption when calculating self-attention, but MIMIR pre-training discards 75% patches, making it still efficient.
Table 9: Pre-training hyperparameters.
@@ Appendix C Details of Training Hyperparameters
Table 10: Fine-tuning hyperparameters.
@@
layer-wise lr decay 0.65, 1.0
@@
warmup epochs 10, 20
training epochs 100, 300
@@
Table 11: Different learning rates. Fine-tuned for 50 epochs.
Dataset Models LR Natural PGD10
CIFAR-10 ViT-T 5.0e-4 76.30 47.60
1.0e-3 80.69 49.56
1.0e-2 85.62 48.78
5.0e-2 85.12 50.30
1.0e-1 84.51 50.40
Figure 4: Natural and adversarial accuracy of ViT-S and ViT-B that are fine-tuned with different numbers of pre-train epochs under a 20-step PGD attack.
Appendix D Training Epochs Evaluation Study
Pre-training epoch. Our experiments so far are based on 800-epoch pre-training. In Figure 4, we show the influence of different numbers of epochs. We use ViT-T and ViT-S as the models for CIFAR-10, ViT-S, and ViT-B for Tiny-ImageNet. We fine-tune each model 50 epochs after MIMIR pre-training. Both adversarial and natural accuracy are improved with a longer training schedule. On CIFAR-10, the improvement of ViT-T is more significant than that of ViT-S. On Tiny-ImageNet, the increase is more obvious. Therefore, a longer pre-training schedule obviously increases performance.
Fine-tuning epoch. To show that the performance of the final model is largely attributed to MIMIR pre-training, we do a very short fine-tuning for pre-trained models. This is to train the randomly initialized classification layer since we do not have the classification layer at pre-training. Therefore, we evaluate the pre-training performance. In Figure 5, we show that MIMIR pre-training plus 5 or 10 epochs of fine-tuning is enough to achieve similar performance compared to 100-epoch fine-tuning.
Figure 5: Natural and adversarial accuracy of ViT-S adversarially fine-tuned for 5 or 10 epochs on CIFAR-10, Tiny-ImageNet, and ImageNet-1K.
Table 12: The average time consumption of a single epoch on 4 GPUs. The “mem.” refers to GPU memory usage.
CIFAR-10 Tiny-ImageNet ImageNet-1K
Architecture #Parames (M) Method time[S] mem.[GB] time[S] mem.[GB] time[S] mem.[GB]
ViT-S 21.34 PGD10 AT 149.3 2.54$\times$4 306.0 3.99$\times$4 2251.7 12.5$\times$4
FastAT 43.3 2.54$\times$4 67.7 4.03$\times$4 555.5 10.4$\times$4
MAE 16.1 3.24$\times$4 33.0 3.27$\times$4 269.6 11.1$\times$4
MIMIR 18.4 3.12$\times$4 40.0 3.18$\times$4 275.5 11.1$\times$4
ViT-B 85.27 PGD10 AT 361.2 5.39$\times$4 1022.2 8.30$\times$4 5416.7 22.1$\times$4
FastAT 122.8 5.36$\times$4 180.2 8.34$\times$4 1361.3 19.8$\times$4
MAE 53.0 5.95$\times$4 106.5 5.95$\times$4 490.9 17.0$\times$4
MIMIR 59.0 6.08$\times$4 122.0 6.11$\times$4 509.9 17.0$\times$4
ConViT-S 27.05 PGD10 AT 442.5 6.64$\times$4 897.0 12.19$\times$4 6626.5 32.5$\times$4
FastAT 106.5 5.86$\times$4 183.2 10.62$\times$4 1431.2 26.4$\times$4
MAE 33.0 10.6$\times$4 67.5 10.61$\times$4 609.7 27.5$\times$4
MIMIR 45.0 10.4$\times$4 90.0 10.54$\times$4 611.1 28.3$\times$4
Appendix E Efficiency
Table 12 provides the time consumption and memory usage of MIMIR and adversarial training at fine-tuning. The time consumption and memory usage are evaluated on 4 A6000 GPUs. Compared to 10-step PGD and FastAT, MIMIR shows much less time consumption. At pre-training, MIMIR shows slightly higher time consumption than MAE, resulting from generating adversarial examples and calculating MI between $x+\delta$ and $z$. Still, as we use a one-step algorithm, generating adversarial examples does not cost too much time. We also consider using a 10-step PGD at pre-training. Table 5 shows that using a one-step algorithm with larger $\alpha=10$ provides better performance. In addition, pre-training with MIMIR also costs less memory. Note in Table 12 that the batch size for MIMIR pre-training is much larger than the batch size for adversarial fine-tuning. In Table 12, a single epoch of MIMIR costs about $\frac{1}{8}$ time consumption compared to 10-step PGD adversarial training, and about $\frac{1}{2}$ compared to 1-step fast adversarial training (Wong, Rice, and Kolter 2020). Therefore, MIMIR is more promising regarding training costs when conducting adversarial training.
Appendix F Mutual Information and HSIC